OpenStudy (anonymous):

Please help me identify the vertex and x-intercepts of f(x)=3/5(x^2+6x-5) and then check the results by writing the quadratic function in standard form??

OpenStudy (anonymous):

@phi do you think you could help me with this one? its the same as the one we did earlier, but I can't find any way to factor the equation

OpenStudy (phi):

that's because it does not factor (nicely) are you sure you got the correct equation (x^2+6x PLUS 5 ) does factor) otherwise we have to use the quadratic formula

OpenStudy (anonymous):

no, sadly it is f(x)=3/5(x^2+6x-5)

OpenStudy (phi):

then use the quadratic formula to find the x-intercepts. In other words, you solve for x when \[ \frac{3}{5}\left(x^2+6x-5\right)=0 \] we can multiply both sides by 5/3 to get rid of the 3/5 \[ \frac{5}{3}\cdot \frac{3}{5}\left(x^2+6x-5\right)=0 \cdot \frac{5}{3} \\ x^2+6x-5=0 \]

OpenStudy (anonymous):

I got x^2+6x=5 x(x+6)=5

OpenStudy (anonymous):

I think that it would be x=0 and x=-6 but im not sure what to do with the 5 on the other side

OpenStudy (phi):

sadly, with a quadratic equation you can't solve it that way (as you just noticed) Have you learned either "completing the square" or the quadratic formula ?

OpenStudy (anonymous):

i've learned both ways but im not very good at them :o

OpenStudy (phi):

let's try completing the square. you got to x^2+6x=5 now instead of factoring out an x, you "complete the square". find the number in front of the x term. divide that number by 2. square it. add the result to both sides of the equation.

OpenStudy (anonymous):

but there is no nummber in front of the x term anymore right?

OpenStudy (phi):

in other words, find the 6 in front of 6x. divide by 2: 6/2=3. square 3 to get 3*3=9

OpenStudy (anonymous):

oh i see

OpenStudy (isaiah.feynman):

 @ali1029